# Hidden Twin Star Solutions from an Agnostic Speed-of-Sound Model: Confronting XTE J1814–338’s Extreme Compactness

### v1

@@ 1 Introduction
Thanks to the high-precision X-ray timing data provided by NASA’s Neutron Star Interior Composistion Explorer (NICER) (Gendreau et al., 2016), recent detailed X-ray pulse modeling of the accreting millisecond X-ray pulsar XTE J1814–338 (Kini et al., 2024) has revealed that this object exhibits an unusually small mass and radius compared to previous MR measurements. Some studies have proposed that this anomaly could be explained by a hybrid star configuration, a neutron star featuring an inner core of exotic matter (e.g., deconfined quark matter) surrounded by a conventional hadronic envelope (Laskos-Patkos & Moustakidis, 2024; Veselsky et al., 2024), or even a dark matter admixed neutron star (Lopes & Issifu, 2025). In this study, we adopt an agnostic construction–widely used in NICER posterior analyses–to first explore the potential of the CS model to produce twin star configurations, and then compare the predicted twin star MR posterior with the anomalous observation of XTE J1814–338 (J1814).
@@ 2.4 Inference Dataset
Finally, we consider the recent observation of XTE J1814–338, analyzed by Kini et al. (2024), The MR posterior file we are using here is the Bkg constrained case. This source is of particular interest due to its unusually small mass and radius, which may indicate the presence of a twin star configuration and is consistent with the compact branch we are investigating.
@@ 3.3 Direct Constraint on Compact Branch from XTE J1814–338
From this inference, we can extract the central density of J1814 based on this given EoS model. For the constraint implemented by only J1814 constraint on compact branch, the posterior central density of J1814 is $\log_{10}(\varepsilon/\mathrm{g\ cm^{-3}})=15.855^{+0.058}_{-0.069}$ and for constraint introduced by whole NICER observations plus J1814. we have the posterior central density equals to $\log_{10}(\varepsilon/\mathrm{g\ cm^{-3}})=15.895^{+0.035}_{-0.042}$ This is indeed an ultra-heavy configuration, with very large central density.
@@ Acknowledgements
H.C. extends gratitude to Mark Alford, Alexander Chen and Ryan O’Connor for insightful discussion.

### v2

@@ 1 Introduction
Recently, (Gendreau et al., 2016), detailed X-ray pulse modeling of the accreting millisecond X-ray pulsar XTE J1814–338 using RXTE data (Kini et al., 2024) has revealed that this object exhibits an unusually small mass and radius compared to previous MR measurements. Some studies have proposed that this anomaly could be explained by a hybrid star configuration, a neutron star featuring an inner core of exotic matter (e.g., deconfined quark matter) surrounded by a conventional hadronic envelope (Laskos-Patkos & Moustakidis, 2024; Veselsky et al., 2024), or even a dark matter admixed neutron star (Lopes & Issifu, 2025; Pitz & Schaffner-Bielich, 2025). In this study, we adopt an agnostic construction–widely used in NICER posterior analyses–to first explore the potential of the CS model to produce twin star configurations, and then compare the predicted twin star MR posterior with the anomalous observation of XTE J1814–338 (J1814).
@@ 2.4 Inference Dataset
Finally, we consider the RXTE observation of XTE J1814–338, recently analyzed by Kini et al. (2024), The MR posterior file we are using here is the Bkg constrained case. This source is of particular interest due to its unusually small mass and radius, which may indicate the presence of a twin star configuration and is consistent with the compact branch we are investigating.
@@ 3.3 Direct Constraint on Compact Branch from XTE J1814–338
From this inference, we can extract the central density of J1814 based on this given EoS model. For the constraint implemented by only J1814 constraint on compact branch, the posterior central density of J1814 is $\log_{10}(\varepsilon/\mathrm{g\ cm^{-3}})=15.855^{+0.058}_{-0.069}$ and for constraint introduced by whole NICER observations plus J1814. We have the posterior central density equals to $\log_{10}(\varepsilon/\mathrm{g\ cm^{-3}})=15.895^{+0.035}_{-0.042}$ This is indeed an ultra-dense configuration, with very large central density.
@@ Acknowledgements
H.C. extends gratitude to Mark Alford, Alexander Chen and Ryan O’Connor for insightful discussion. All the reproduction files and necessary datasets are available in Zenodo (Zhou & Huang, 2025).